OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4)

hero (hero):

1. Find the slope, m 2. input one of the points (x,y) along with m into the formula y = mx+b 3. solve for b 4. re-write the formula y = mx + b and insert the values you found for m and b into it.

OpenStudy (anonymous):

m=3/2 and then like... y=3/2(1)+b and then I get stuck... :(

hero (hero):

To substitute m = 3/2 and point (1,1): y = mx + b 1 = 3/2(1) + b

hero (hero):

What did you get for b?

OpenStudy (anonymous):

would b be -1/2?

OpenStudy (anonymous):

-1/2+3/2=1?

hero (hero):

yes

hero (hero):

So what do you get for the final result?

OpenStudy (anonymous):

I think it's -1/2+3/2=1 maybe?

hero (hero):

The final form has to be y = mx + b, with only values for m and b included.

OpenStudy (anonymous):

idk...D:

hero (hero):

y = 3/2x - 1/2 Notice that only values for m and b have been included. x and y stay remain as variables.